OpenStudy (anonymous):

write the equation of a line that includes the points (-4,0) and a slope of (-3/5) in slope intercept form

OpenStudy (pinklion23):

ok so what lesson number is this

OpenStudy (anonymous):

idk its on gradpoint

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

After that solve for y

OpenStudy (anonymous):

umm i don't know how to do it

pooja195 (pooja195):

Ok you are given the slope (m) m=-3/5 x1=-4 y1=0 Plug it into this formula (Point slope formula) \[\huge~\rm~y-0=-3/5(x-(-4)\] \[\huge~\rm~y-0=-3/5(x+4)\] Now solve for y

OpenStudy (anonymous):

that -3/5 is actaully a fraction

pooja195 (pooja195):

Ye its your slope.

pooja195 (pooja195):

In the formula m=slope

OpenStudy (anonymous):

um can i just have the answer please cause im still not understand it

pooja195 (pooja195):

If you don't try you won't get it in the future I suggest you try.

OpenStudy (anonymous):

i have tried my notebooks right beside me

OpenStudy (anonymous):

give me a sec and i can help

OpenStudy (anonymous):

is it y=-3/5x+12/5 ?

pooja195 (pooja195):

y=-3/5x-12/5 So yes you are correct but its negative... as i corrected above

OpenStudy (anonymous):

thansk

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)